# second generation 3G third generation

### v1

@@ IV-1 TALS-TUCKER receiver
    For the TALS-TUCKER receiver, the problems (47), (48), and (49) are unique in the LS sense if $\boldsymbol{V}_{1}=[\boldsymbol{\mathcal{C}}]_{(1)}\big(\overline{\boldsymbol{G}}\otimes\boldsymbol{\Psi}\otimes\boldsymbol{X}\big)^{\text{T}}\in\mathbb{C}^{N\times IKT}$, $\boldsymbol{V}_{2}=\boldsymbol{\mathcal{C}}]_{(2)}\big(\overline{\boldsymbol{G}}\otimes\boldsymbol{\Psi}\otimes\boldsymbol{HS}\big)^{\text{T}}\in\mathbb{C}^{M_{T}\times IKM_{R}}$ and $\boldsymbol{V}_{4}=[\boldsymbol{\mathcal{C}}]_{(4)}\big(\boldsymbol{\Psi}\otimes\boldsymbol{X}\otimes\boldsymbol{HS}\big)^{\text{T}}\in\mathbb{C}^{M_{T}N\times KTM_{R}}$ are full row rank. Satisfying these conditions implies the following inequalities
@@ IV-2 PAKRON receiver
    The conditions necessary to ensure a unique joint channel and symbol recovery for the PAKRON receiver are associated with stage I. In this sense, the problems (27) and (28) are unique in the LS sense if $\boldsymbol{R}_{1}=\overline{\boldsymbol{G}}\diamond\boldsymbol{\Psi}\in\mathbb{C}^{IK\times M_{T}N}$ and $\boldsymbol{R}_{2}=\boldsymbol{\Psi}\diamond\boldsymbol{\Omega}\in\mathbb{C}^{KTM_{R}\times M_{T}N}$ are full column rank, which implies the following conditions to be satisfied:
@@
    Additionally, the stage I of the PAKRON receiver enables the derivation of a sufficient uniqueness condition. Since $\boldsymbol{\mathcal{Z}}$ follows a third-order PARAFAC model, uniqueness is ensured if the so-called Kruskal condition is satisfied [stegeman2007]
    $Kr(\boldsymbol{\Omega})+Kr(\boldsymbol{\Psi})+Kr(\overline{\boldsymbol{G}})\geq 2M_{T}N+2.$ (58)
    However, the received signal tensor exploited by the PAKRON receiver involves structured matrices. More specifically, in the PARAFAC decomposition (22), the first and second mode factor matrices have Kronecker and Khatri-Rao product structures, respectively, as shown in (18) and (19). Therefore, it is important to analyze how these structures influence the uniqueness condition (58). To this end, note that $\textrm{rank}(\boldsymbol{\Omega})=\textrm{rank}(\boldsymbol{X}\otimes\boldsymbol{HS})=\textrm{rank}(\boldsymbol{X})\textrm{rank}(\boldsymbol{HS})$. Considering $T\geq M_{T}$, and using the fact the $\boldsymbol{S}\in\mathbb{C}^{N\times N}$ is full rank by construction, we have
    $\textrm{rank}(\boldsymbol{\Omega})=T\textrm{rank}(\boldsymbol{H})=T\min(M_{R},N).$ (59)
    Resorting to the lemma on the rank of the Khatri-Rao product (see [SidNway2000, STEGEMAN2007540, LDL2008]), and using the fact that $\boldsymbol{W}$ and $\boldsymbol{P}$ are full-rank matrices by construction, $\boldsymbol{\Psi}$ has full-rank equal to $K$. Additionally, assuming scattering-rich propagation the Kruskal-rank condition (58) can be rewritten as a function of the system parameters as follows
@@
    It should be noted from (60) that various scenarios can be explored based on different system parameter configurations. In this context, it is also worth noting that the PAKRON receiver still operates in a more challenging scenario under rank-deficient channel models, as it may not satisfy condition (60). For instance, in the extreme case where the rank of the RIS-BS channel $\boldsymbol{H}$ is one (which arises under line of sight (LoS) propagation between the BS and the RIS), we have from (59) that $\textrm{rank}(\boldsymbol{\Omega})=T$. In this case, when we collect a sufficient number of frames so that $I\geq M_{T}N$, condition (60) is automatically fulfilled. Otherwise, if $I<M_{T}N$, it becomes necessary to have $T+K+I\geq 2M_{T}N+2$ in order to compensate for the rank deficiency of the channels.
@@
    Complexity: In algorithms involving pseudo-inverse calculations or rank-$1$ approximations, the computational complexity is generally dominated by the Singular Value Decomposition (SVD) [Favier2019]. Note that, for an arbitrary matrix $\boldsymbol{A}\in\mathbb{C}^{M\times N}$, the complexity of its SVD computation is $\boldsymbol{\mathcal{O}}(MN\min(M,N))$. It can be noted that the complexity of solving an LS problem is dominated by the associated SVD computations. Considering that the PAKRON algorithm comprises two stages, we have: In Stage I, the computational complexity is primarily determined by steps $1$ and $2$, which have complexities of $\boldsymbol{\mathcal{O}}((M_{T}N)^{2}IK)$ and $\boldsymbol{\mathcal{O}}((M_{T}N)^{2}KTM_{R})$, respectively. In Stage II, the dominant computation corresponds to a single rank-$1$ approximation, which has complexity $\boldsymbol{\mathcal{O}}(M_{T}NTM_{R})$. Then, the asymptotic complexity per iteration of the PAKRON receiver is $\max\big((M_{T}N)^{2}IK,(M_{T}N)^{2}KTM_{R},M_{T}NTM_{R}\big)$. On the other hand, the computational complexity of the TALS-TUCKER receiver is dictated by steps $1$, 22, and $3$, which have complexities $\boldsymbol{\mathcal{O}}((N)^{2}IKT)$, $\boldsymbol{\mathcal{O}}(IKM_{R}(M_{T})^{2})$, and $\boldsymbol{\mathcal{O}}(KTM_{R}(M_{T}N)^{2})$, respectively. Therefore, the asymptotical complexity of the TALS-TUCKER receiver is $\max\big((N)^{2}IKT,IKM_{R}(M_{T})^{2},KTM_{R}(M_{T}N)^{2}\big)$. Table I summarizes the asymptotic complexities the identifiability conditions of our proposed receivers in comparison to competing methods. It is important to emphasize that BALS ([Sokal_BD_RIS_2024]) is the simplest receiver and, consequently, generally exhibits the lowest computational complexity in terms of flops when compared with the BTKF, PAKRON, and TUCKER receivers. Moreover, as the number of BD-RIS scattering elements increases, the TUCKER receiver becomes less complex than the BTKF one, which by its turn, is less complex than PAKRON (although this gap noticeably narrows as $N$ grows). Among all considered schemes, PAKRON usually incurs the highest flop count. Note, however, that the higher computational complexity of our proposed methods comes with their ability to operate in semi-blindly without the need of pilot sequences.
@@ V Simulation results
    $\textrm{NMSE}(\hat{\mathbf{H}})={\color[rgb]{0,0,0}\frac{1}{R}\sum_{r=1}^{R}\dfrac{\|\boldsymbol{\Xi}^{(r)}-\hat{\boldsymbol{\Xi}}^{(r)}\|_{F}^{2}}{\|\boldsymbol{\Xi}^{(r)}\|_{F}^{2}}},$ (61)
    where $\boldsymbol{\Xi}\in\{\boldsymbol{H},\overline{\boldsymbol{G}}\}$ and $\hat{\boldsymbol{\Xi}}^{(r)}$ is their estimate obtained in $r$-th Monte Carlo run. We also show numerical results in terms of the symbol error rate (SER) performance as a function of the signal to noise ratio (SNR) in dB. As references for comparison, we also plot the performance of competing (pilot-assisted) baseline methods, such as the LS-based channel estimation method proposed in [Clerckx_CAMSAP_2023], and the BTKF and BTALS methods proposed recently in [Sokal_BD_RIS_2024]. The performance of the traditional zero-forcing receiver operating under perfect channel knowledge is also plotted. The results represent an average over $R=1000$ Monte Carlo runs. Regarding the channel model, we consider the Rayleigh fading case (i.e., the entries of the channel matrices are independent and identically distributed zero-mean circularly symmetric complex Gaussian random variables) as well as a multipath (low-rank) geometrical channel model, where the path directions are randomly generated according to a uniform distribution. In this case, at each Monte Carlo run, the azimuth and elevation angles are drawn within the intervals $\pi/2,\pi/2]$ and $0,\pi/2]$, respectively. The system parameters used in this analysis were $N=16$, $M_{R}=4$, $M_{T}=2$, $I=2$, $K=32$, $T=4$, and $Q\in\{2,4,8\}$.
@@
    Figures 6 and 7 illustrate the channel estimation accuracy PAKRON and TUCKER receivers. Let us first consider Figure 6, which depicts the performance of the estimated user to the BR-RIS channel ($\overline{\boldsymbol{G}}$). It is noteworthy that the PAKRON receiver shows greater degradation than the TUCKER receiver at lower SNR levels (in particular until 1010dB). However, in the higher SNR regime, this performance gap tends to diminish, resulting in nearly identical performances, with the PAKRON receiver showing a slight tendency to outperform the TUCKER one. Moreover, we observe that both algorithms are sensitive to variations in the number of groups $Q$. This sensitivity is anticipated since the number of blocks can influence the iterative part of the PAKRON receiver, specifically the estimation of the channel $\overline{\boldsymbol{G}}$ and the structured matrix $\boldsymbol{\Omega}$. On the other hand, Figure 7 demonstrates that the TUCKER receiver consistently outperforms the PAKRON receiver across the entire SNR range. Nevertheless, both solutions show insensitivity to the number of groups. This can be explained by: $(I)$ For the PAKRON receiver, the estimate of the channel $\boldsymbol{H}$ is extracted by applying a Kronecker factorization to the $\boldsymbol{\Omega}$ (stage II) matrix already estimated during the iterative part (stage I). This process reduces the sensitivity of the $\boldsymbol{H}$ channel estimation concerning variations in the number of groups. However, there is error propagation from stage I to stage II, which explains the gap between the PAKRON and TUCKER receivers, thereby favoring the latter. In summary, PAKRON operates on a third-order tensor model whose factor matrices are constrained. This induces a coupling between the dimensions of the original tensor. As a consequence, it becomes more sensitive to changes in the number of groups and suffers a performance degradation in the estimation of channel $\boldsymbol{G}$ at low SNR when compared with the TUCKER receiver. The TUCKER receiver, in contrast, preserves the fourth-order tensor structure of the received signal and thus does not incur diversity loss, since no dimension coupling occurs. As a result, it achieves more accurate estimates of $\boldsymbol{G}$ in the low SNR regime. Nevertheless, this advantage tends to diminish as $Q$ grows. For the channel $\boldsymbol{H}$, the TUCKER receiver also surpasses the PAKRON receiver in all simulated SNR range. This is mainly because it retains the diversity offered by the fourth-order tensor structure while enabling a single-stage processing, thereby avoiding the extra stage required by PAKRON. Such an additional stage leads to error propagation, which appears to be the primary reason for the performance disparity between the two receivers with respect to the estimation of the channel matrix $\boldsymbol{H}$.
@@
    In Figure 11, we present a compelling comparison of our algorithms against established works in the literature, specifically the innovative proposal in [B_Clerckc_CE_TSP_24] and the advanced approach outlined in [Sokal_Asilomar_2024]. The author of [B_Clerckc_CE_TSP_24] provides a conventional LS solution, while [Sokal_Asilomar_2024] and [Sokal_BD_RIS_2024] introduce a sophisticated tensor approach, culminating in a closed-form solution known as LS-KRF[Sokal_Asilomar_2024], BTKF [Sokal_BD_RIS_2024], and BTALS [Sokal_BD_RIS_2024]. for composite channel estimation. This one is characterized as a block Kronecker product between the involved channels, expressed as: $\boldsymbol{H}_{C}=\boldsymbol{H}|\otimes|\boldsymbol{G}=[\boldsymbol{H}^{(1)}|\otimes|\boldsymbol{G}^{(1)}\dots\boldsymbol{H}^{(Q)}|\otimes|\boldsymbol{G}^{(Q)}]$. Unlike our approach, which assumes the channel $\boldsymbol{G}$ varies dynamically over time, the competitor’s model assumes a quasi-static channel. To ensure a fair and rigorous comparison, the NMSE of our composite channel is calculated as an average [^6] An alternative approach could involve considering $I=1$, but this scenario falls within the framework addressed in our conference version across $I$ composite channels, where $\boldsymbol{H}_{C}^{i}=\boldsymbol{H}|\otimes|\boldsymbol{G}_{i}$ for $i=1,\dots,I$. Considering the setup $M_{R}=10$, $M_{T}=2$, $K=32$, $I=4$, $T=8$, $N=16$ and an i.i.d channel model, our proposed solutions, the TUCKER and PAKRON receivers, demonstrate good performance, both significantly outperforming the LS method described in [B_Clerckc_CE_TSP_24]. This outcome is expected, given that the LS solution’s simplicity fails to leverage the intrinsic tensor structure of the signal model. In contrast, our approach leverages this structure to achieve noise-rejection gain. Moreover, while the competitor receiver analyzed in [Sokal_Asilomar_2024] and [Sokal_BD_RIS_2024] deliver superior results, this is understandable since both our method and the competitor algorithms benefit from effective noise rejection. However, our semi-blind receivers necessitate the estimation of three matrices ($\boldsymbol{H}$, $\overline{\boldsymbol{G}}$, and $\boldsymbol{X}$), compared to LS-KRF, BTALS, and BTKF receivers which are limited to estimating the first two. Nevertheless, the proposed semi-blind methods also delivers early estimates of the data symbols, in contrast to the competing methods which do not have this feature, requiring a dedicated pilot transmission period prior to symbol decoding.
@@
    The outcomes are influenced by the designs of matrices $\boldsymbol{P}$ and $\boldsymbol{W}$, and consequently by $\boldsymbol{\Psi}$. We focus on the design outlined in Section II-A. It is essential to note that the optimal design of the rotation matrix $\mathbf{P}$ and the coding matrix $\mathbf{W}$ can significantly impact the results. However, exploring this aspect is beyond the scope of this paper.

### v2

@@ IV-1 TALS-TUCKER receiver
    For the TALS-TUCKER receiver, the problems (47), (48), and (49) are unique in the LS sense if $\bm{V}_{1}=[\bm{\mathcal{C}}]_{(1)}\big(\overline{\bm{G}}\otimes\bm{\Psi}\otimes\bm{X}\big)^{\text{T}}\in\mathbb{C}^{N\times IKT}$, $\bm{V}_{2}=\bm{\mathcal{C}}]_{(2)}\big(\overline{\bm{G}}\otimes\bm{\Psi}\otimes\bm{HS}\big)^{\text{T}}\in\mathbb{C}^{M_{T}\times IKM_{R}}$ and $\bm{V}_{4}=[\bm{\mathcal{C}}]_{(4)}\big(\bm{\Psi}\otimes\bm{X}\otimes\bm{HS}\big)^{\text{T}}\in\mathbb{C}^{M_{T}N\times KTM_{R}}$ are full row rank. Satisfying these conditions implies the following inequalities
@@ IV-2 PAKRON receiver
    The conditions necessary to ensure a unique joint channel and symbol recovery for the PAKRON receiver are associated with stage I. In this sense, the problems (27) and (28) are unique in the LS sense if $\bm{R}_{1}=\overline{\bm{G}}\diamond\bm{\Psi}\in\mathbb{C}^{IK\times M_{T}N}$ and $\bm{R}_{2}=\bm{\Psi}\diamond\bm{\Omega}\in\mathbb{C}^{KTM_{R}\times M_{T}N}$ are full column rank, which implies the following conditions to be satisfied:
@@
    Additionally, the stage I of the PAKRON receiver enables the derivation of a sufficient uniqueness condition. Since $\bm{\mathcal{Z}}$ follows a third-order PARAFAC model, uniqueness is ensured if the so-called Kruskal condition is satisfied [stegeman2007]
    $Kr(\bm{\Omega})+Kr(\bm{\Psi})+Kr(\overline{\bm{G}})\geq 2M_{T}N+2.$ (58)
    However, the received signal tensor exploited by the PAKRON receiver involves structured matrices. More specifically, in the PARAFAC decomposition (22), the first and second mode factor matrices have Kronecker and Khatri-Rao product structures, respectively, as shown in (18) and (19). Therefore, it is important to analyze how these structures influence the uniqueness condition (58). To this end, note that $\textrm{rank}(\bm{\Omega})=\textrm{rank}(\bm{X}\otimes\bm{HS})=\textrm{rank}(\bm{X})\textrm{rank}(\bm{HS})$. Considering $T\geq M_{T}$, and using the fact the $\bm{S}\in\mathbb{C}^{N\times N}$ is full rank by construction, we have
    $\textrm{rank}(\bm{\Omega})=T\textrm{rank}(\bm{H})=T\min(M_{R},N).$ (59)
    Resorting to the lemma on the rank of the Khatri-Rao product (see [SidNway2000, STEGEMAN2007540, LDL2008]), and using the fact that $\bm{W}$ and $\bm{P}$ are full-rank matrices by construction, $\bm{\Psi}$ has full-rank equal to $K$. Additionally, assuming scattering-rich propagation the Kruskal-rank condition (58) can be rewritten as a function of the system parameters as follows
@@
    It should be noted from (60) that various scenarios can be explored based on different system parameter configurations. In this context, it is also worth noting that the PAKRON receiver still operates in a more challenging scenario under rank-deficient channel models, as it may not satisfy condition (60). For instance, in the extreme case where the rank of the RIS-BS channel $\bm{H}$ is one (which arises under line of sight (LoS) propagation between the BS and the RIS), we have from (59) that $\textrm{rank}(\bm{\Omega})=T$. In this case, when we collect a sufficient number of frames so that $I\geq M_{T}N$, condition (60) is automatically fulfilled. Otherwise, if $I<M_{T}N$, it becomes necessary to have $T+K+I\geq 2M_{T}N+2$ in order to compensate for the rank deficiency of the channels.
@@
    Complexity: In algorithms involving pseudo-inverse calculations or rank-$1$ approximations, the computational complexity is generally dominated by the Singular Value Decomposition (SVD) [Favier2019]. Note that, for an arbitrary matrix $\bm{A}\in\mathbb{C}^{M\times N}$, the complexity of its SVD computation is $\bm{\mathcal{O}}(MN\min(M,N))$. It can be noted that the complexity of solving an LS problem is dominated by the associated SVD computations. Considering that the PAKRON algorithm comprises two stages, we have: In Stage I, the computational complexity is primarily determined by steps $1$ and $2$, which have complexities of $\bm{\mathcal{O}}((M_{T}N)^{2}IK)$ and $\bm{\mathcal{O}}((M_{T}N)^{2}KTM_{R})$, respectively. In Stage II, the dominant computation corresponds to a single rank-$1$ approximation, which has complexity $\bm{\mathcal{O}}(M_{T}NTM_{R})$. Then, the asymptotic complexity per iteration of the PAKRON receiver is $\max\big((M_{T}N)^{2}IK,(M_{T}N)^{2}KTM_{R},M_{T}NTM_{R}\big)$. On the other hand, the computational complexity of the TALS-TUCKER receiver is dictated by steps $1$, 22, and $3$, which have complexities $\bm{\mathcal{O}}((N)^{2}IKT)$, $\bm{\mathcal{O}}(IKM_{R}(M_{T})^{2})$, and $\bm{\mathcal{O}}(KTM_{R}(M_{T}N)^{2})$, respectively. Therefore, the asymptotical complexity of the TALS-TUCKER receiver is $\max\big((N)^{2}IKT,IKM_{R}(M_{T})^{2},KTM_{R}(M_{T}N)^{2}\big)$. Table I summarizes the asymptotic complexities the identifiability conditions of our proposed receivers in comparison to competing methods. It is important to emphasize that BALS ([Sokal_BD_RIS_2024]) is the simplest receiver and, consequently, generally exhibits the lowest computational complexity in terms of flops when compared with the BTKF, PAKRON, and TUCKER receivers. Moreover, as the number of BD-RIS scattering elements increases, the TUCKER receiver becomes less complex than the BTKF one, which by its turn, is less complex than PAKRON (although this gap noticeably narrows as $N$ grows). Among all considered schemes, PAKRON usually incurs the highest flop count. Note, however, that the higher computational complexity of our proposed methods comes with their ability to operate in semi-blindly without the need of pilot sequences.
@@ V Simulation results
    $\textrm{NMSE}(\hat{\mathbf{H}})={\color[rgb]{0,0,0}\frac{1}{R}\sum_{r=1}^{R}\dfrac{\|\bm{\Xi}^{(r)}-\hat{\bm{\Xi}}^{(r)}\|_{F}^{2}}{\|\bm{\Xi}^{(r)}\|_{F}^{2}}},$ (61)
    where $\bm{\Xi}\in\{\bm{H},\overline{\bm{G}}\}$ and $\hat{\bm{\Xi}}^{(r)}$ is their estimate obtained in $r$-th Monte Carlo run. We also show numerical results in terms of the symbol error rate (SER) performance as a function of the signal to noise ratio (SNR) in dB. As references for comparison, we also plot the performance of competing (pilot-assisted) baseline methods, such as the LS-based channel estimation method proposed in [Clerckx_CAMSAP_2023], and the BTKF and BTALS methods proposed recently in [Sokal_BD_RIS_2024]. The performance of the traditional zero-forcing receiver operating under perfect channel knowledge is also plotted. The results represent an average over $R=1000$ Monte Carlo runs. Regarding the channel model, we consider the Rayleigh fading case (i.e., the entries of the channel matrices are independent and identically distributed zero-mean circularly symmetric complex Gaussian random variables) as well as a multipath (low-rank) geometrical channel model, where the path directions are randomly generated according to a uniform distribution. In this case, at each Monte Carlo run, the azimuth and elevation angles are drawn within the intervals $\pi/2,\pi/2]$ and $0,\pi/2]$, respectively. The system parameters used in this analysis were $N=16$, $M_{R}=4$, $M_{T}=2$, $I=2$, $K=32$, $T=4$, and $Q\in\{2,4,8\}$.
@@
    Figures 6 and 7 illustrate the channel estimation accuracy PAKRON and TUCKER receivers. Let us first consider Figure 6, which depicts the performance of the estimated user to the BR-RIS channel ($\overline{\bm{G}}$). It is noteworthy that the PAKRON receiver shows greater degradation than the TUCKER receiver at lower SNR levels (in particular until 1010dB). However, in the higher SNR regime, this performance gap tends to diminish, resulting in nearly identical performances, with the PAKRON receiver showing a slight tendency to outperform the TUCKER one. Moreover, we observe that both algorithms are sensitive to variations in the number of groups $Q$. This sensitivity is anticipated since the number of blocks can influence the iterative part of the PAKRON receiver, specifically the estimation of the channel $\overline{\bm{G}}$ and the structured matrix $\bm{\Omega}$. On the other hand, Figure 7 demonstrates that the TUCKER receiver consistently outperforms the PAKRON receiver across the entire SNR range. Nevertheless, both solutions show insensitivity to the number of groups. This can be explained by: $(I)$ For the PAKRON receiver, the estimate of the channel $\bm{H}$ is extracted by applying a Kronecker factorization to the $\bm{\Omega}$ (stage II) matrix already estimated during the iterative part (stage I). This process reduces the sensitivity of the $\bm{H}$ channel estimation concerning variations in the number of groups. However, there is error propagation from stage I to stage II, which explains the gap between the PAKRON and TUCKER receivers, thereby favoring the latter. In summary, PAKRON operates on a third-order tensor model whose factor matrices are constrained. This induces a coupling between the dimensions of the original tensor. As a consequence, it becomes more sensitive to changes in the number of groups and suffers a performance degradation in the estimation of channel $\bm{G}$ at low SNR when compared with the TUCKER receiver. The TUCKER receiver, in contrast, preserves the fourth-order tensor structure of the received signal and thus does not incur diversity loss, since no dimension coupling occurs. As a result, it achieves more accurate estimates of $\bm{G}$ in the low SNR regime. Nevertheless, this advantage tends to diminish as $Q$ grows. For the channel $\bm{H}$, the TUCKER receiver also surpasses the PAKRON receiver in all simulated SNR range. This is mainly because it retains the diversity offered by the fourth-order tensor structure while enabling a single-stage processing, thereby avoiding the extra stage required by PAKRON. Such an additional stage leads to error propagation, which appears to be the primary reason for the performance disparity between the two receivers with respect to the estimation of the channel matrix $\bm{H}$.
@@
    In Figure 11, we present a compelling comparison of our algorithms against established works in the literature, specifically the innovative proposal in [B_Clerckc_CE_TSP_24] and the advanced approach outlined in [Sokal_Asilomar_2024]. The author of [B_Clerckc_CE_TSP_24] provides a conventional LS solution, while [Sokal_Asilomar_2024] and [Sokal_BD_RIS_2024] introduce a sophisticated tensor approach, culminating in a closed-form solution known as LS-KRF[Sokal_Asilomar_2024], BTKF [Sokal_BD_RIS_2024], and BTALS [Sokal_BD_RIS_2024]. for composite channel estimation. This one is characterized as a block Kronecker product between the involved channels, expressed as: $\bm{H}_{C}=\bm{H}|\otimes|\bm{G}=[\bm{H}^{(1)}|\otimes|\bm{G}^{(1)}\dots\bm{H}^{(Q)}|\otimes|\bm{G}^{(Q)}]$. Unlike our approach, which assumes the channel $\bm{G}$ varies dynamically over time, the competitor’s model assumes a quasi-static channel. To ensure a fair and rigorous comparison, the NMSE of our composite channel is calculated as an average [^6] An alternative approach could involve considering $I=1$, but this scenario falls within the framework addressed in our conference version across $I$ composite channels, where $\bm{H}_{C}^{i}=\bm{H}|\otimes|\bm{G}_{i}$ for $i=1,\dots,I$. Considering the setup $M_{R}=10$, $M_{T}=2$, $K=32$, $I=4$, $T=8$, $N=16$ and an i.i.d channel model, our proposed solutions, the TUCKER and PAKRON receivers, demonstrate good performance, both significantly outperforming the LS method described in [B_Clerckc_CE_TSP_24]. This outcome is expected, given that the LS solution’s simplicity fails to leverage the intrinsic tensor structure of the signal model. In contrast, our approach leverages this structure to achieve noise-rejection gain. Moreover, while the competitor receiver analyzed in [Sokal_Asilomar_2024] and [Sokal_BD_RIS_2024] deliver superior results, this is understandable since both our method and the competitor algorithms benefit from effective noise rejection. However, our semi-blind receivers necessitate the estimation of three matrices ($\bm{H}$, $\overline{\bm{G}}$, and $\bm{X}$), compared to LS-KRF, BTALS, and BTKF receivers which are limited to estimating the first two. Nevertheless, the proposed semi-blind methods also delivers early estimates of the data symbols, in contrast to the competing methods which do not have this feature, requiring a dedicated pilot transmission period prior to symbol decoding.
@@
    The outcomes are influenced by the designs of matrices $\bm{P}$ and $\bm{W}$, and consequently by $\bm{\Psi}$. We focus on the design outlined in Section II-A. It is essential to note that the optimal design of the rotation matrix $\mathbf{P}$ and the coding matrix $\mathbf{W}$ can significantly impact the results. However, exploring this aspect is beyond the scope of this paper.